OpenStudy (anonymous):

the half life of radium 226 is 1590 years. if a sample initially contains 100mg, how many milligrams will remain after 1 week

OpenStudy (anonymous):

PHYSICS !!!

OpenStudy (whpalmer4):

If the half-life of Ra-226 is 1590 years, then the amount remaining at time \(t\) (in years) will be the initial amount multiplied by the following:\[\frac{1}{2^{t/1590}}\]or \[2^{-t/1590}\] After 1 half life, (1590 years), the multiplier is \(2^{-1590/1590} = 2^{-1} = \frac{1}{2}\) Now you need to figure out what fraction of a year is 1 week...or retool the equation to use 1590 years in weeks.